solve for x log4log3log2(x3-7x2+14x)

solve for x


log4log3log2(x3-7x2+14x) 

Grade:12

1 Answers

Aman Bansal
13 years ago

dear Abdul Yazdan Qadri

log4log3log2(x3 -7x2 +14x) =0

log3log2(x3 -7x2 +14x)=40

log2 (x3 -7x2 +14x)=31

(x-1)(x-2)(x-4)=0

hence x=1,2or 4


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
AMAN BANSAL